$SPAD/src/input f02aff.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
ia:=4
n:=4
a:Matrix SF:=
 [[  1.5 ,0.1 , 4.5 ,-1.5 ],_
  [-22.5 ,3.5 ,12.5 ,-2.5 ],_
  [ -2.5 ,0.3 , 4.5 ,-2.5 ],_
  [ -2.5 ,0.1 , 4.5 , 2.5 ]]
result:=f02aff(ia,n,a,-1)

———-

References^{[1] nothing}

4